# Ultrastable nanophotonic microcavities via integrated thermometry

### v1

@@ I Thermal Design for Integrated Thermometry
$\rho C_{p}\frac{\partial T}{\partial t}+\nabla\cdot\bm{q}=Q+Q_{ted}$ (M1)
where $\rho$ is the material density, $C_{p}$ is the specific heat capacity at constant stress, $T$ represents the temperature, $q=-k\nabla T$ is the heat flux by conduction, $k$ is the thermal conductivity, $Q$ contains additional heat sources, and $Q_{ted}$ accounts for thermo-elastic damping in the material. All material thermal properties were set by the standard COMSOL material database for the cross-section of our devices. Figure M1(a) displays the cross-section of our device, and the steady-state spatial temperature profile of the stationary thermal simulation. The heat power is set to $100$ mW for the purposes of illustration. The initial temperature of the entire chip, and the boundary condition temperature for the bottom face of the chip, are set to $293$ K. Here, we denote the resulting steady-state temperature change at the center of the SiN core and at the center of the resistance thermometer as $\delta$Tcore and $\delta$Ttherm., respectively.
Figure M1(b) shows the temperature differential $\delta$Tcore - $\delta$Ttherm. (left y-axis, in blue) as well as the heater power required to heat up the core (right y-axis, in dark red), which is a measure of heater efficiency, versus the heater-thermometer spacing. For small spacing, the heater-thermometer and heater-core separations are considerably different, leading to a large temperature differential. However, the close proximity allows for greater heater efficiency. For large spacing, the temperature differential drops, as the two separations are almost equal, while the heater efficiency worsens. The design employed in this work utilizes a $20$ $\mu$m heater-thermometer spacing, which leads to a very low temperature differential of $<5$ mK, making it effective in stabilizing the resonance frequency of the microcavity. The design principles described here can be readily applied to other integrated material platforms.
II Device Fabrication
III Measurement of Temperature Coefficient of Resistance (TCR)
Figure M2: Measurement of temperature coefficient of resistance of the integrated Platinum thermometer. a, Experimental schematic depicting a test chip, containing a platinum resistor (referred to here as device under test (DUT)) fabricated on top of a SiN microresonator, placed on a temperature-stabilized hot plate chuck inside a vacuum chamber. Its electrical resistance is probed as a function of temperature by performing current-voltage characterization using a Keithley $2400$ series SourceMeter®. b, Measured electrical resistance of DUT versus temperature (dark-blue filled circles) plotted alongside a linear fit (magenta solid line).
@@ III Measurement of Temperature Coefficient of Resistance (TCR)
The integrated Platinum resistance thermometer, which allows the real-time measurement of the microresonator as described in the Main Text, is calibrated as shown in Fig. M2(a). A test device is placed on a temperature-controlled chuck using a thermally-conductive adhesive in order to ensure proper thermal contact. The setup is placed and operated inside a vacuum probe station in order to minimize unwanted convective cooling. Two probes, connected to a Keithley $2400$ series SourceMeter®, are brought in electrical contact with the two ends of the Pt resistor under test (labeled as device under test (DUT) in Fig. M2). At a given temperature of the chuck, the resistance $R$ of the DUT is measured by performing a current-voltage characterization. The voltage drop across the two contact pads is measured at various source currents, and a linear fit of the resulting data is obtained to extract $R$. This method is further explained in a prior manuscript [52]. The source current values used are low enough ($<1$ mA) to avoid any self-heating of the resistor. This process is repeated for a range of chuck temperatures $T$. Figure M2(b) displays the results. Over the range of temperatures probed on this experiment, the resistance exhibits a linear trend. The temperature-coefficient of resistance is extracted by fitting the experimental data to Eq. 1, and is determined to be $A=0.00172$. We note that DUT described here has a shorter fabricated length compared to the resistance thermometer described in the Main Text, leading to the lower base resistance.

### v2

@@ III Thermal Design for Integrated Thermometry
$\rho C_{p}\frac{\partial T}{\partial t}+\nabla\cdot\bm{q}=Q+Q_{ted}$ (E1)
[…]
IV Measurement of Temperature Coefficient of Resistance (TCR)
@@
The integrated Platinum resistance thermometer, which allows the real-time measurement of the microresonator as described in the Main Text, is calibrated as shown in Fig. 2(a). A test device is placed on a temperature-controlled chuck using a thermally-conductive adhesive in order to ensure proper thermal contact. The setup is placed and operated inside a vacuum probe station in order to minimize unwanted convective cooling. Two probes, connected to a Keithley $2400$ series SourceMeter®, are brought in electrical contact with the two ends of the Pt resistor under test (labeled as device under test (DUT) in Fig. 2). At a given temperature of the chuck, the resistance $R$ of the DUT is measured by performing a current-voltage characterization. The voltage drop across the two contact pads is measured at various source currents, and a linear fit of the resulting data is obtained to extract $R$. This method is further explained in a prior manuscript [52]. The source current values used are low enough (<1<1 mA) to avoid any self-heating of the resistor. This process is repeated for a range of chuck temperatures $T$. Figure 2(b) displays the results. Over the range of temperatures probed on this experiment, the resistance exhibits a linear trend. The temperature-coefficient of resistance is extracted by fitting the experimental data to Eq. E2, and is determined to be $A=0.00172$. We note that DUT described here has a shorter fabricated length compared to the resistance thermometer described in the Main Text, leading to the lower base resistance.
V Measurement of Waveguide Thermo-Optic Coefficient
A small-amplitude 2-Hz sinusoidal voltage is applied to the heater while monitoring the electrical resistance of the thermometer. Simultaneously, the shift in resonance frequency is probed using a scanning laser centered at 1560 nm, as described above. The free-spectral-range (FSR) of the SiN microresonator is 76 GHz, its loaded $Q$ is approximately $3\times 10^{6}$, and the resonances in the telecommunications C-band are critically coupled. The width of the waveguide is 2100 nm. The measured change in thermometer resistance is converted to a change in temperature using the following equation,
$R(T)=R_{0}(1+A\Delta T)$ (E2)
where $R_{0}$ is the resistance at room temperature, $\Delta T$ is the change in temperature, and $A$ is the temperature coefficient of resistance (TCR). The coefficient $A$ is measured independently to be $A=0.00172$. The measurement procedure is detailed in Extended Data Fig. 2.
Extended Data Figure 1: Thermal design for integrated thermometry. a, Numerically simulated steady-state temperature profile for the scheme described in this work. The heater power is set to $100$ mW for the purposes of illustration. The initial temperature of the chip is $293$ K. b, Heater-thermometer spacing design trade-off, showing the core-thermometer temperature differential (left y-axis, in blue), and the heater power required for core heating of $0.1$ K (right y-axis, in dark red), versus heater-thermometer spacing.
Extended Data Figure 2: Measurement of temperature coefficient of resistance of the integrated Platinum thermometer. a, Experimental schematic depicting a test chip, containing a platinum resistor (referred to here as device under test (DUT)) fabricated on top of a SiN microresonator, placed on a temperature-stabilized hot plate chuck inside a vacuum chamber. Its electrical resistance is probed as a function of temperature by performing current-voltage characterization using a Keithley $2400$ series SourceMeter®. b, Measured electrical resistance of DUT versus temperature (dark-blue filled circles) plotted alongside a linear fit (magenta solid line).
VI Data Availability
The data shown in this paper are available in this Zenodo repository: https://doi.org/10.5281/zenodo.17144471.